# ICon: In-Context Contribution for Automatic Data Selection

### v1

@@ Abstract
Data selection for instruction tuning is essential for improving the performance of Large Language Models (LLMs) and reducing training cost. However, existing automated selection methods either depend on computationally expensive gradient-based measures or manually designed heuristics, which may fail to fully exploit the intrinsic attributes of data. In this paper, we propose In-context Learning for Contribution Measurement (ICon), a novel gradient-free method that takes advantage of the implicit fine-tuning nature of in-context learning (ICL) to measure sample contribution without gradient computation or manual indicators engineering. ICon offers a computationally efficient alternative to gradient-based methods and reduces human inductive bias inherent in heuristic-based approaches. ICon comprises three components and identifies high-contribution data by assessing performance shifts under implicit learning through ICL. Extensive experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets demonstrate the effectiveness of ICon. Remarkably, on LLaMA3.1-8B, models trained on 15% of ICon-selected data outperform full datasets by 5.42% points and exceed the best performance of widely used selection methods by 2.06% points. We further analyze high-contribution samples selected by ICon, which show both diverse tasks and appropriate difficulty levels, rather than just the hardest ones. [^1] Project page: https://annayang2020.github.io/ICon_Data_Selection/
ICon: In-Context Contribution for Automatic Data Selection
[…]
@@ 1 Introduction
Figure 1: Overview of the ICon Method. The diagram illustrates the two key components of the method: Contribution Quantification with the ICon Score and ICon-guided Selection Paradigm Training. This approach enables gradient-free, bias-reduced data selection with scalable inference.
@@
Recent advancements in large language models (LLMs), such as GPT-4o Hurst et al. (2024), have demonstrated strong capabilities in both understanding and generation. These models are capable of performing a wide range of tasks Liu et al. (2023b); Dong et al. (2023); Team et al. (2024); Guo et al. (2025), often showing impressive flexibility and adaptability. Instruction tuning Mishra et al. (2021); Wei et al. (2021); Sanh et al. (2022) has emerged as a powerful approach to enhance the performance of such models by finetuning them to better follow human instructions. While traditional instruction tuning relies on amassing vast datasets Köpf et al. (2023); Chung et al. (2024), recent study Zhou et al. (2023a) manually selecting data has shown that training with carefully selected subsets of high-quality data can achieve superior performance with reduced computational cost.
Automated data selection methods are essential, as manually selecting the most valuable training data is time-consuming and impractical. Previous studies on automated data selection Xia et al. (2024); Zhang et al. (2024) have explored gradient-based approaches to estimate sample value, but these methods are often computationally expensive and inefficient. Other works propose lightweight alternatives based on manually designed textual features Liu et al. (2023c); Bukharin and Zhao (2023); Zhang et al. (2025) or simple heuristics such as difficulty Li et al. (2023a, 2024b) and uncertainty Kung et al. (2023). However, these methods fail to directly measure the training impact of samples and introduce human inductive bias into the selection process. Specifically, manually defined indicators, such as difficulty, uncertainty, or topic coverage, can not accurately capture the true contribution of a sample, and reliance on predefined metrics limits the ability to fully exploit the potential of the training data.
In this paper, we introduce In-Context Learning for Contribution Measurement (ICon), a novel selection method that captures the contribution of individual samples while avoiding human inductive bias and costly gradient computation. Inspired by in-context learning (ICL), our method takes advantadge of recent findings Dong et al. (2022); Dai et al. (2023); Zheng et al. (2023); Yin et al. (2024) that ICL simulates parameter updates via the attention mechanism, effectively acting as implicit fine-tuning without the need for explicit gradient updates. Our key intuition is that this implicit finetuning nature allows us to directly assess the contribution of individual samples, without the need for manually defined indictor engineering. Additionally, since ICon does not require explicit training or gradient computation, it enables highly efficient selection, significantly reducing computational overhead compared to traditional gradient-based methods.
@@
Our approach leverages ICL to measure sample contribution through three components: 1) an assessment set to provide a reliable performance reference, 2) ICon scores that quantify contribution via ICL’s implicit fine-tuning, and 3) a lightweight selection paradigm for scalable data selection. The ICon score captures each sample’s contribution without costly gradient computation or human-designed heuristics. A higher positive score indicates a greater contribution, while a lower negative score reflects more detrimental effects. Finally, the ICon-guided selection paradigm is trained efficiently with LoRA Hu et al. (2021) based on the target model, reducing the selection complexity to linear inference calls.
@@
We demonstrate the effectiveness of ICon across multiple models and evaluation settings. Experiments on LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024) are conducted using 12 widely used benchmarks and 5 pairwise comparison test sets. Models trained on a small fraction of ICon-selected data outperform their full-data counterparts. Notably, with only 15% of the data, the LLaMA3.1-8B and Qwen2.5-3B models improve by 5.42 and 1.24 percentage points, respectively; LLaMA2-7B achieves a 0.65-point gain using just 5%. ICon also surpasses widely used selection methods. For example, ICon improves LLaMA3.1-8B’s average benchmark score by 2.06 points over the best prior method. Furthermore, we also analyze the impact of data scale, cross-dataset generalization, and the properties of high-contribution samples selected by ICon. The main contributions of our paper are as follows:
  • We propose ICon, a novel data selection method inspired by ICL. By leveraging its implicit fine-tuning nature, ICon enables gradient-free and bias-reduced measurement of each sample’s contribution to model performance.
  • Our approach avoids gradient updates and auxiliary models, reducing computational cost and design complexity. Once trained with LoRA, the ICon-guided selection paradigm enables data selection through linear-complexity inference calls, making it highly scalable.
@@
  • We demonstrate the effectiveness of ICon through experiments on multiple models, 12 widely used benchmarks, and 5 pairwise test sets, showing consistent performance gains over baselines.
@@ 2 Related Work
Data Selection
Data selection aims to construct optimal datasets to improve model performance Murphy (2012), reduce training cost Suárez et al. (2019); Sorscher et al. (2022), mitigate undesirable model behaviors Longpre et al. (2024), and ensure evaluation quality Oren et al. (2024). With the rise of LLMs, it plays a central role across various training stages, including pretraining, instruction tuning, alignment, in-context learning, and task-specific finetuning. Pretraining selection focuses on filtering large-scale raw data Soldaini et al. (2024); Penedo et al. (2023); Instruction tuning often uses manual Zhou et al. (2023a) or automated methods, with the latter detailed in the next section; alignment involves manual curation Bai et al. (2022), model-based evaluation Wang et al. (2023b); Cui et al. (2023), and reward model re-weighting Touvron et al. (2023); Pace et al. (2024); in-context learning emphasizes demonstration choice Xu and Zhang (2024); Luo et al. (2024) and order Lu et al. (2021); and task-specific tuning leverages utility-based Ivison et al. (2023) or empirical methods Grangier and Iter (2022).
Data Selection for Instruction Tuning
Instruction tuning Wei et al. (2021); Liu et al. (2023a); Xu et al. (2024) is a common approach to enhance instruction-following capabilities. Recent work Zhou et al. (2023a) shows that carefully selecting a smaller, higher quality dataset can improve both the effectiveness and efficiency of instruction tuning. Existing automatic data selection methods fall into three main categories: (1) selection based on textual features like helpfulness, quality, or complexity Zhang et al. (2025); Cao et al. (2023); Chen et al. (2023); Liu et al. (2023c); Bukharin and Zhao (2023); Minghao et al. (2024); (2) selection using training-time gradients to capture model-specific characteristics Xia et al. (2024); Zhang et al. (2024); Pan et al. (2024); Chang et al. (2024); (3) selection guided by feedback during inference, which is model-aware and much more efficient Du et al. (2023); Kung et al. (2023); Li et al. (2023a, 2024b); Liu et al. (2024); Wang et al. (2024); Li et al. (2024a). The first and third approaches rely on handcrafted heuristics for scoring and filtering, introducing human inductive bias, while the second demands substantial computational cost. ICon addresses the limitations of existing approaches by avoiding both costly gradient computations and manually designed heuristics. It leverages ICL to directly measure sample contribution in a model-aware, efficient manner.
@@ 3 Methodology
To efficiently assess the contribution of each sample, we draw on recent advances in in-context learning (ICL), which has been recognized as a form of implicit finetuning without explicit gradient updates Dai et al. (2023); Yin et al. (2024). Specifically, ICL leverages the model’s attention mechanism to simulate parameter updates by processing demonstration examples, thereby adapting the model’s behavior within the given context. The key intuition is that in-context learning can effectively simulate the training impact of individual samples, enabling an assessment of their influence on model performance without the high computational cost of finetuning.
Building on this insight, we propose the ICon method, an efficient data selection method that harnesses ICL’s implicit finetuning nature to assess each sample’s contribution to overall model performance. As shown in Figure 1, ICon comprises three phases: (1) constructing a representative assessment set, (2) computing the proposed ICon score to quantify sample contribution, and (3) learning an ICon-guided selection paradigm to curate data for final instruction tuning.
@@ 3.1 Assessment Set Construction
The goal of this phase is to construct an assessment set that reliably reflects the model’s overall capabilities. To ensure diversity and reduce bias, we sample data from three sources: ChatGPT-generated, GPT-4-generated, and human-authored data, resulting in a high-quality set covering a range of task types and instruction complexities. The assessment set is used solely for assessing model performance, with no overlap with training data or downstream benchmarks, ensuring independent evaluation and fair comparison. Formally, we define the assessment set as $D_{a}$, consisting of $n$ instruction tuning task samples, where each is represented as a pair $(x^{a},y^{a})$. Here, $x=\text{map}(\text{Instruction},[\text{Input}])$ represents the full instruction, which is formed by concatenating the instruction and optional input, and $y$ denotes the corresponding response. Similarly, we also define the candidate training set as $D_{t}=\{(x^{t}_{1},y^{t}_{1}),(x^{t}_{2},y^{t}_{2}),\ldots,(x^{t}_{m},y^{t}_{m})\}$, where $m$ is the number of candidate training instruction-response pairs.
3.2 Contribution Quantification with ICon Score
In this phase, we introduce the ICon score, a metric for estimating the contribution of individual samples to overall model performance. The core idea is to leverage ICL to evaluate contribution of sample. ICL enables models to adapt to new information presented within a prompt, providing an efficient approach to assess the contribution of individual samples on task performance. We adopt perplexity as the assessment metric, as it directly evaluates the likelihood of response tokens and serves as a smoother indicator of model performance compared to accuracy Li et al. (2023a, 2024b); Jiang et al. (2024). For a task sample $S_{i}=(x^{a}_{i},y^{a}_{i})\in D_{a}$, where $D_{a}$ is the assessment set, $y^{a}_{i}$ contains $N$ tokens and the $k$-th token is denoted as $y^{a}_{i,k}$, the model’s performance under parameters $\theta$ on task $i$ is defined as:
@@ 3.2 Contribution Quantification with ICon Score
Then, We select a sample $T_{j}=(x^{t}_{j},y^{t}_{j})$ from $D_{t}^{\prime}$, a subset of the training set $D_{t}$. This sample is inserted into the context during inference to assess the model’s performance on an assessment sample $S_{i}$ after implicitly learning from $T_{j}$. The model’s performance on $S_{i}$, given $T_{j}$ in context is denoted as $\text{PPL}_{\theta}(S_{i}|T_{j})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j},x^{a}_{i})$. Since perplexity is affected by both content and input length, we introduce a fairness adjustment to neutralize length effects. Specifically, we define the baseline performance on $S_{i}$ without the influence of $T_{j}$ by using a randomly generated meaningless sequence $T_{j}^{\text{rand}}$, which matches the length of $T_{j}$. The model’s baseline performance on $T_{i}$ is then denoted as $\text{PPL}_{\theta}(S_{i}|T_{j}^{\text{rand}})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j}^{\text{rand}},x^{a}_{i})$.
Using these values, we define the task-level ICon score of sample $S_{j}$ on task $i$ under parameters $\theta$ as shown in Equation 2, where $\epsilon$ is a small constant to avoid division by zero. A positive task-level ICon score indicates that $S_{j}$ contributes positively to model performance on task $i$, while a negative score suggests a detrimental effect. We then compute the global ICon score by aggregating the task-level scores across all assessment tasks, as shown in Equation 3. This global score allows us to rank samples based on their overall contribution. A higher positive global ICon score implies a stronger positive impact on model overall performance, while a lower negative score indicates the opposite.
@@
Figure 2: Pairwise comparison results (win-tie-lose) between models trained on ICon-selected subsets and models trained on the full Alpaca dataset across five evaluation benchmarks: Vicuna, Koala, WizardLM, SInstruct, and LIMA. ICon-selected data enhances instruction-following ability with fewer samples, as evidenced by pairwise comparisons.
3.3 ICon-guided Selection Paradigm Training
While the ICon score effectively captures a sample’s contribution, it requires multiple inference calls per candidate training sample. To further improve efficiency, we train an ICon-guided selection paradigm using the computed global ICon scores. This allows us to reduce the selection complexity to $O(m)$ inference calls, where $m$ is the total number of candidate samples. Specifically, we use a subset of the training data and label the top $K\%$ of samples by global ICon score as high-contribution. A selection model is then trained using LoRA on the target model to classify high-value samples. This paradigm is applied to the full candidate set to select the most contributive samples for instruction tuning, thereby enhancing the model’s ability to follow instructions.

### v2

# RICo: Refined In-Context Contribution for Automatic Instruction-Tuning Data Selection

@@ Abstract
Data selection for instruction tuning is crucial for improving the performance of large language models (LLMs) while reducing training costs. In this paper, we propose Refined Contribution Measurement with In-Context Learning (RICo), a novel gradient-free method that quantifies the fine-grained contribution of individual samples to both task-level and global-level model performance. RICo enables more accurate identification of high-contribution data, leading to better instruction tuning. We further introduce a lightweight selection paradigm trained on RICo scores, enabling scalable data selection with a strictly linear inference complexity. Extensive experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets demonstrate the effectiveness of RICo. Remarkably, on LLaMA3.1-8B, models trained on 15% of RICo-selected data outperform full datasets by 5.42% points and exceed the best performance of widely used selection methods by 2.06% points. We further analyze high-contribution samples selected by RICo, which show both diverse tasks and appropriate difficulty levels, rather than just the hardest ones. [^1] https://annayang2020.github.io/ICon_Data_Selection/
RICo: Refined In-Context Contribution for Automatic
Instruction-Tuning Data Selection
Yixin Yang, Qingxiu Dong, Linli Yao, Fangwei Zhu, Zhifang Sui State Key Laboratory of Multimedia Information Processing, Peking University {yangyx,dqx,linliyao,zhufangwei2022}@stu.pku.edu.cn, szf@pku.edu.cn
@@ 1 Introduction
Figure 1: Overview of the RICo Method. The diagram illustrates the two key components of the method: refined contribution quantification with the RICo score and RICo-guided selection paradigm training. This approach enables gradient-free, bias-reduced data selection with scalable inference.
@@
Recent advancements in large language models (LLMs), such as GPT-4o Hurst et al. (2024), have demonstrated strong capabilities in both understanding and generation. These models are capable of performing a wide range of tasks Liu et al. (2023a); Dong et al. (2023); Team et al. (2024); Guo et al. (2025), often showing impressive flexibility and adaptability. Instruction tuning Mishra et al. (2021); Wei et al. (2021); Sanh et al. (2022); Liu et al. (2023b); Xu et al. (2024) has emerged as a powerful approach to enhance the performance of such models by finetuning them to better follow human instructions. While traditional instruction tuning relies on amassing vast datasets Köpf et al. (2023); Chung et al. (2024), recent work Zhou et al. (2023a) shows that manual selection of high-quality data subsets can achieve better performance with lower computational cost.
Automated data selection is crucial, as manual selection is costly and impractical. Previous studies on automated data selection Xia et al. (2024); Zhang et al. (2024a) have explored gradient-based approaches to estimate sample value, but these are often computationally expensive. Others propose lightweight methods based on manually designed heuristics, such as textual features Liu et al. (2023c); Bukharin and Zhao (2023); Zhang et al. (2025) and instruction difficulty Li et al. (2023a, 2024b), but these do not directly assess training impact and can introduce human inductive bias into the selection process. Recent works Li et al. (2023c); Jiao et al. (2025) adopt binary comparisons in in-context learning (ICL) settings. While the implicit fine-tuning nature of ICL Dong et al. (2022); Dai et al. (2023); Zheng et al. (2023); Yin et al. (2024) shows promise in addressing prior limitations, existing approaches remain limited in robustly assessing sample value. They provide only coarse-grained signals, are often biased toward longer sample Wang et al. (2022a), and typically require a large number of inference calls.
In this paper, we introduce Refined Contribution Measurement with In-Context Learning (RICo), a novel gredient-free selection method that captures the fine-grained contribution of individual samples to both task-level performance and overall model capability. RICo measures each sample’s contribution across diverse tasks while reducing biases such as length sensitivity and human inductive bias. We argue that such refined measurement enables more accurate identification of high-contribution data, leading to stronger instruction tuning results. Moreover, we train a lightweight selection paradigm on these contributions, reducing inference calls and enabling scalable data selection over large candidate pools.
@@
Our approach measures the refined contribution of each sample through three components: 1) an assessment set that provides a reliable reference for model performance, 2) RICo scores that quantify the fine-grained contribution, and 3) a lightweight selection paradigm for scalable data selection. The RICo score reflects the contribution of each sample to task-specific and overall performance, including a fairness adjustment to mitigate length-related bias. A higher positive score indicates a greater contribution, while a lower negative score reflects a more detrimental effect. Finally, the RICo-guided selection paradigm is efficiently trained with LoRA Hu et al. (2021) on the target model, reducing selection complexity to linear inference calls with respect to the number of training samples.
@@
We demonstrate the effectiveness of RICo across multiple models and evaluation settings. Experiments on LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024) are conducted using 12 widely used benchmarks and 5 pairwise comparison test sets. Models trained on a small fraction of RICo-selected data outperform their full-data counterparts. Notably, with only 15% of the data, the LLaMA3.1-8B and Qwen2.5-3B models improve by 5.42 and 1.24 percentage points, respectively; LLaMA2-7B achieves a 0.65-point gain using just 5%. RICo also surpasses widely used selection methods. For example, RICo improves LLaMA3.1-8B’s average benchmark score by 2.06 points over the best prior method. Furthermore, we also analyze the impact of data scale, cross-dataset generalization, and the properties of high-contribution samples selected by RICo. The main contributions of our paper are as follows:
  • We propose RICo, a novel gradient-free data selection method that quantifies the refined contribution of individual samples to both task-level and overall model performance, with fairness adjustments to reduce length-related bias.
  • We introduce a lightweight selection paradigm trained on RICo contribution scores, enabling scalable data selection with a strictly linear inference complexity.
@@
  • We demonstrate the effectiveness of RICo through experiments on multiple models, 12 widely used benchmarks, and 5 pairwise test sets, showing consistent performance gains over baselines.
@@ 2 Related Work
Automatic Data Selection
Automatic data selection seeks to algorithmically construct optimal datasets to improve model performance Murphy (2012), reduce training cost Suárez et al. (2019); Sorscher et al. (2022), mitigate undesirable model behaviors Longpre et al. (2024), and ensure evaluation quality Oren et al. (2024). With the rise of LLMs, it plays a central role across various training stages, including pretraining, instruction tuning, alignment, in-context learning, and task-specific finetuning Albalak et al. (2024). Pretraining selection focuses on filtering large-scale raw data Soldaini et al. (2024); Penedo et al. (2023); instruction tuning will be discussed in the next section; alignment involves model-based evaluation Wang et al. (2023b); Cui et al. (2023) and reward model re-weighting Touvron et al. (2023); Pace et al. (2024); in-context learning emphasizes demonstration choice Xu and Zhang (2024); Luo et al. (2024) and order Lu et al. (2021); and task-specific tuning leverages utility-based Ivison et al. (2023) or empirical methods Grangier and Iter (2022).
Automatic Instruction-Tuning Data Selection
Instruction-tuning data selection aims to identify the most informative subsets of data to improve model performance during instruction tuning. Gradient-based methods Xia et al. (2024); Zhang et al. (2024a); Pan et al. (2024); Chang et al. (2024) assess sample influence through gradients or subset training outcomes, but incur substantial computational cost. Heuristic-based methods Liu et al. (2023c); Bukharin and Zhao (2023); Liu et al. (2024); Li et al. (2024b); Zhang et al. (2025) use manually crafted indicators, which may not fully capture the true training impact and can introduce human inductive bias. Appendix D contains the full list of heuristic-based methods cited in this work. In-context probing (ICP) methods Li et al. (2023c); Jiao et al. (2025) leverage the implicit fine-tuning effect of ICL without model updates, reducing both human bias and computational cost. However, existing approaches typically rely on simple binary comparisons, yielding coarse-grained signals, exhibiting length bias, and incurring inference costs that grow multiplicatively with the size of the candidate pool. Inspired by the advantages of ICL, RICo introduces a fine-grained, bias-reduced contribution measurement and a lightweight selection paradigm that enables scalable data selection with strictly linear inference complexity.
@@ 3 Methodology
To accurately and efficiently identify high-contribution samples from a training candidate set, we propose RICo, a novel gradient-free data selection method that measures the refined contribution of individual samples. As illustrated in Figure 1, RICo consists of three components: 1) constructing a representative assessment set, 2) computing the RICo score to quantify fine-grained sample contribution, and 3) training a RICo-guided selection model to curate the full dataset with strictly linear complexity.
@@ 3.1 Assessment Set Construction
The goal of this component is to construct an assessment set that spans diverse tasks and reliably reflects the model’s overall capabilities. To promote reliability and mitigate bias in data origin and task design, we sample from three sources: ChatGPT-generated, GPT-4-generated, and human-authored data, resulting in a high-quality set covering various task types and instruction complexities. The assessment set is used solely for assessment, with no overlap with training data or downstream benchmarks, ensuring independence and fair comparison. Formally, we define the assessment set as $D_{a}$, consisting of $n$ instruction tuning samples, each represented as a pair $(x^{a},y^{a})$, where $x^{a}=\text{map}(\text{Instruction},[\text{Input}])$ is the full instruction formed by concatenating the instruction and the optional input, and $y^{a}$ is the corresponding response. Meanwhile, the candidate training set is denoted as $D_{t}=\{(x^{t}_{1},y^{t}_{1}),(x^{t}_{2},y^{t}_{2}),\ldots,(x^{t}_{m},y^{t}_{m})\}$, where $m$ is the number of instruction-response pairs.
3.2 Refined Contribution Quantification
We introduce the RICo score, a principled metric for quantifying the refined contribution of individual samples to model performance. It enables fine-grained assessment at both task-specific and global levels, while being low-cost and bias-reduced.
Perplexity
Perplexity directly evaluates the likelihood of response tokens and provides a smoother measure of model performance compared to accuracy. It aligns well with instruction tuning objectives, making it a commonly adopted metric for both evaluation and data selection in instruction-tuning tasks Li et al. (2023a, 2024b); Jiang et al. (2024); Li et al. (2023c). Specifically, for a task sample $S_{i}=(x^{a}_{i},y^{a}_{i})\in D_{a}$, where $D_{a}$ is the assessment set, $y^{a}_{i}$ consists of $N$ tokens and the $k$-th token is denoted as $y^{a}_{i,k}$, the model’s performance under parameters $\theta$ on task $i$ is defined as shown in Equation LABEL:eq:example1. To evaluate the contribution of a training sample $T_{j}=(x^{t}_{j},y^{t}_{j})$ from a subset $D_{t}^{\prime}\subseteq D_{t}$, the sample is inserted into the context during inference. The model’s performance on $S_{i}$ is then reassessed after implicitly conditioning on $T_{j}$. The perplexity in this setting is denoted as Equation 2.
@@ 3.2 Refined Contribution Quantification
$\text{PPL}_{\theta}(S_{i}\mid T_{j})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j},x^{a}_{i})$ (2)
Fairness Adjustment
Despite its wide use, we observe that perplexity tends to decrease with longer inputs, even when the content is uninformative Wang et al. (2022a). This length sensitivity introduces a bias toward longer samples and fails to fairly reflect true sample contribution. To address this, we introduce a fairness adjustment to neutralize length effects. Specifically, we define the baseline performance on $S_{i}$ by replacing $T_{j}$ with a randomly generated, semantically meaningless sequence $T_{j}^{\text{rand}}$ of the same length. The length-controlled reference perplexity is defined as:
$\text{PPL}_{\theta}(S_{i}\mid T_{j}^{\text{rand}})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j}^{\text{rand}},x^{a}_{i})$ (3)
Task-level RICo Score
We further define the task-level RICo score to quantify the refined contribution of an individual sample on model behavior for a specific task. The score integrates the above fairness adjustment to reduce length bias. We also normalize for task difficulty to reduce its interference in contribution scoring, ensuring that the score highlights the effect of the sample rather than the task’s inherent difficulty. Specifically, we define the task-level RICo score of sample $S_{j}$ on task $i$ under parameters $\theta$ as shown in Equation 4, where $\epsilon$ is a small constant to avoid division by zero. A higher task-level RICo score indicates a greater contribution of $S_{j}$ to model performance on task $i$.
@@
Global-level RICo Score
To quantify a sample’s overall contribution to the model’s performance, we compute the global-level RICo score by aggregating its task-level scores. The diversity and coverage of the assessment set provide a reliable reference for reflecting overall model performance. As we aim to improve generalization rather than specialization in any particular task type, we assign equal weight to each task. The global-level score, shown in Equation 5, reflects the refined overall contribution of a sample and enables ranking based on total impact. Higher global-level RICo scores indicate greater overall contribution to model performance.
Figure 2: Pairwise comparison results (win-tie-lose) between models trained on RICo-selected subsets and models trained on the full Alpaca dataset across five evaluation benchmarks: Vicuna, Koala, WizardLM, SInstruct, and LIMA. RICo-selected data enhances instruction-following ability with fewer samples, as evidenced by pairwise comparisons.
3.3 RICo-guided Selection Paradigm
To further improve selection efficiency, we train an RICo-guided selection model using the computed global-level RICo scores. This reduces the selection complexity from $O(nm)$ to $O(m)$ inference calls, where $n$ is the number of assessment samples and $m$ is the number of training candidates. As a result, it enables efficient and scalable identification of high-contribution samples for instruction tuning. Specifically, we sample a subset of training data and label the top $K\%$ of samples by global-level RICo score as high-contribution. A selection model is then trained using LoRA on the target model to classify high-contribution samples. Once trained, this selection paradigm can be applied to the full candidate pool, enabling linear-time selection in large-scale settings, with potential for practical use in realistic applications.